OpenStudy (anonymous):

.Simplify square root parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis divided by cosine squared theta. square root sine theta ±sin Θ ±cos Θ |tan Θ|

OpenStudy (jtvatsim):

\[\sqrt{\frac{(1-\cos\theta)(1+\cos\theta)}{\cos^2 \theta}}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Excuse my ugliness.

OpenStudy (jtvatsim):

LOL, no worries. :)

OpenStudy (anonymous):

:3

OpenStudy (jtvatsim):

Well, this problem certainly looks frightening doesn't it?

OpenStudy (anonymous):

yes :c

OpenStudy (jtvatsim):

But, math is known for trying to make itself scarier than it really is. It's just that most people run before they see that math just tries to act scary.

OpenStudy (anonymous):

I Dont like math ;c

OpenStudy (jtvatsim):

I guess most people don't... you are in the majority. :)

OpenStudy (jtvatsim):

neways, lets focus in on one piece at a time. I'm going to write THETA as t if you don't mind...

OpenStudy (anonymous):

okay c:

OpenStudy (jtvatsim):

Alright, so let's take a big breath and look at only one part of the problem... how about tackling those parenthesis first...

OpenStudy (jtvatsim):

Let's deal with (1 - cos t)(1 + cos t) can we simplify this?

OpenStudy (jtvatsim):

Or expand?

OpenStudy (anonymous):

Yes. o-o would it be sin?

OpenStudy (jtvatsim):

Hmm... not sure, how did you get that?

OpenStudy (anonymous):

i was using pythagoras identity ;o

OpenStudy (jtvatsim):

So, we get (1 - cos t)(1 + cos t) = 1 - cos^2 t right?

OpenStudy (anonymous):

Yess.

OpenStudy (jtvatsim):

Alright, so are you saying that 1 - cos^2 t = sin t ?

OpenStudy (anonymous):

Sin^2t ;o

OpenStudy (jtvatsim):

Good! Now I agree! ;)

OpenStudy (anonymous):

OH WAIT. Would the answer be a tan? ;o

OpenStudy (jtvatsim):

AHA! You've got it on the run! Yes! :D

OpenStudy (anonymous):

So D?

OpenStudy (jtvatsim):

Yes, the absolute value bars are there because we have tan^2 on the inside

OpenStudy (anonymous):

THANK YOUUUU.

OpenStudy (jtvatsim):

YOU ARE WELCOME!! And you did all the thinking for the most part! :)

OpenStudy (jtvatsim):

You did well for not liking math!

OpenStudy (anonymous):

C: